OpenStudy (anonymous):

Expand the following using the properties of logarithms and simplify. Assume when necessary that all quantities represent positive real numbers. ln((sqaure root of (Z))/(xy))

OpenStudy (campbell_st):

well start by rewriting the square root in index form \[\frac{z^{\frac{1}{2}}}{xy}\] now for division, subtract the logs \[\ln(\frac{a}{b}) = \ln(a) - \ln(b)\] the rule for powers is \[\ln(x^a) = a \times \ln(x)\] and lastly the rule for multiplication \[\ln(ab) = \ln(a) + \ln(b)\] you will use all 3 rules to expand you log expression.

OpenStudy (campbell_st):

oops it should be \[\ln(\frac{z^{\frac{1}{2}}}{xy})\]

OpenStudy (anonymous):

so this is what i did previously: (1/2)log(z)-log(x)+log(y) but apparently it was wrong.That is a math problems website where it tells you if you entered any wrong answers. so the answer I entered was wrong.

OpenStudy (campbell_st):

nearly... the log(y) is also negative 1/2 log(z) - log(xy) = 1/2 log(z) - (log(x) + log(y))

OpenStudy (campbell_st):

so you can distribute the negative to get 1/2 log(z) - log(x) - log(y)

OpenStudy (anonymous):

i tried that too but it is also wrong and I don't know what is the problem :/

OpenStudy (campbell_st):

well is it is \[\log(\sqrt{\frac{z}{xy}})\] if it is... then its 1/2 of each log... I worked on \[\log(\frac{\sqrt z}{xy})\] which are different expressions.

OpenStudy (anonymous):

The expression I gave first which you worked on is the correct one.

OpenStudy (campbell_st):

ok... there is a floating bracket in the denominator are you sure its not \[\frac{\log(\sqrt{z})}{xy}\]

OpenStudy (anonymous):

I just got it :3 I kept entering log instead of ln and that was the problem. Thanks for you help any way :)

OpenStudy (campbell_st):

glad to help